OpenStudy (anonymous):

The Jones family attends a movie for $47.00 with two adults and three kids. The Smith family attends when some friends are in town for $74.75 with three adults and five kids. How much does the Ramirez family have to pay with adults and four kids

OpenStudy (anonymous):

set up some equations

OpenStudy (anonymous):

what kind??

OpenStudy (anonymous):

Jones family: 2*adults + 3*kids = $47 total 2*a + 3*k = 47 2a + 3k = 47

OpenStudy (anonymous):

all three of those equations represent the same thing

OpenStudy (anonymous):

can you set up an equation for the Smith family now?

OpenStudy (anonymous):

3a+5k=74.75 whats next?? @bi

OpenStudy (anonymous):

perfect now you have two equations and two unknowns, can you solve this system for a and k? 2a + 3k = 47 3a + 5k = 74.75

OpenStudy (anonymous):

how do you do that? sorry

OpenStudy (anonymous):

are you allowed to use a calculator on tests?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

hmm ok, the easiest way would be either elimination or substitution, which do you prefer?

OpenStudy (anonymous):

substitution

OpenStudy (anonymous):

what is the last sentence supposed to say? "How much does the Ramirez family have to pay with adults and four kids"

OpenStudy (anonymous):

its missing the number of adults

OpenStudy (anonymous):

with four adults and four kids

OpenStudy (anonymous):

a. y=45+3k?

OpenStudy (anonymous):

ok solve one of those 2 equations for either a or k, you choose

OpenStudy (anonymous):

hmm close but not quite

OpenStudy (anonymous):

2a + 3k = 47, lets solve for a, so first subtract 3k from both sides

OpenStudy (anonymous):

2a = 47 - 3k, now divide both sides by 2

OpenStudy (anonymous):

22

OpenStudy (anonymous):

no, a = (47 - 3k) / 2

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

and what will be the answer

OpenStudy (anonymous):

to the whole problem?

OpenStudy (anonymous):

just to the first one

OpenStudy (anonymous):

we are working on it

OpenStudy (anonymous):

do you understand how we got a = (47 - 3k) / 2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so now you know the value of a, we cannot plug it back into the same equation because we will not get any new information, so lets plug it into the other equation

OpenStudy (anonymous):

can you please plug what we found for a, into the other equation?

OpenStudy (anonymous):

3a=74.75-5k

OpenStudy (anonymous):

thats true but you need to plug what we found for a into that equation

OpenStudy (anonymous):

so that we can solve for k

OpenStudy (anonymous):

can u plug it in.. I'm kinda lost

OpenStudy (anonymous):

ok lets start again, because me doing it for you wont help anyone

OpenStudy (anonymous):

we have these two equations that we came up with 2a + 3k = 47 3a + 5k = 74.75

OpenStudy (anonymous):

we looked at the first equation and solved for a 2a + 3k = 47 2a = 47 - 3k a = (47 - 3k) / 2

OpenStudy (anonymous):

now we plug that value for a, into the second equation 3a + 5k = 74.75 3(plug in here) + 5k = 74.75

OpenStudy (anonymous):

go for it

OpenStudy (anonymous):

3(47-3k)+5k=74.75

OpenStudy (anonymous):

thats very close but you left out the divide by 2, i will put it in for you 3((47-3k)/2)+5k=74.75

OpenStudy (anonymous):

$$3*\frac{47-3k}{2} + 5k = 74.75$$

OpenStudy (anonymous):

now you need to solve that for k

OpenStudy (anonymous):

by the way elimination method would have been much easier, but thats ok this is good practice

OpenStudy (anonymous):

so solve that for k and what do you get for k?

OpenStudy (anonymous):

im not sure how you do it

OpenStudy (anonymous):

maybe it will help to write (47 - 3k) / 2 as separate fractions, can you do that?